Karen works in retail computer store. She receives a weekly base salary of $300 plus commission of 3% of sales exceeding her quota of $20,000 per week. What are her sales for a week in which she earns $630.38?
Answer by checkley71(8403)

.03(X-20,000)+300=630.38
.03X-600+300=630.38
.03X=630.38+300
.03X=930.38
X=930.38/.03
X=$31,012.67 IS HER TOTAL SALES.
PROOF
.03(31,012.67-20,000)+300=630.38
.03*11,012.67+300=630.67
330.38+300=630.38
630.38=630.38